OpenStudy (anonymous):

cube root 4a/cube root a^2

OpenStudy (anonymous):

\[\frac{\sqrt[3]{4a}}{\sqrt[3]{a^2}}=\sqrt[3]{4}*a^{1/3}*a^{-2/3}=\sqrt[3]{4}*a^{-1/3}=\sqrt[3]{\frac{4}{a}}\]